An automobile is traveling at 60.0 km/h. Its tires have a radius of 33.0 cm..

a) Find the angular speed of the tires ( in rad/s ).

b) Find the angular displacement of the tires in 30.0 s.

c) Find the linear distance traveled by a point on the tread in 30.0 s.

d) Find the linear distance traveled by the automobile in 30.0 s.

JUST ANSWER C and D. What is the difference between them?
Answer #1

given

v = 60 km/h = 60*5/18 m/s = 16.7 m/s
r = 33 cm = 0.33 m

a) angular speed , w = v/r

= 16.7/0.33

= 50.6 rad/s

b) angular displacement during 30 s,

theta = w*t

= 50.6*30

= 1518 rad (or) 241.6 revolutions

c) the linear distance traveled by a point on the tread in 30.0 s. = r*theta

= 0.33*1518

= 501 m

d) the linear distance traveled by the automobile in 30.0 s. = v*t

= 16.7*30

= 501 m